The perimeter of a rectangular fram is 72 inches. The width is 4 more than the length what are the measurements of the width and the length?
Found 2 solutions by mathsmiles, solver91311:
Answer by mathsmiles(68)

The perimeter of a rectangular frame is W + W + L + L. Or 2W + 2L.

You were given a couple important pieces of information for this problem.
First, P = 72 so
2W + 2L = 72
And you know the W = L + 4.

If we start with the first equation, 2W + 2L = 72 and substitute for W using the 2nd equation, we get:
2(L+4) + 2L = 72 Multiplying out the paren
2L + 8 + 2L = 72 Combining terms:
4L + 8 = 72 Subtracting 8 from each side of the equation:
4L = 64 dividing each side by 4:
L = 16

If the width is 4 more than the length, then
W = 16 + 4 = 20

Checking:
2(16) + 2(20) = 72
32 + 40 = 72
72 = 72 Correct!


Answer by solver91311(24713)



If a rectangle with perimeter has a length that is units longer (or shorter if ) than times the width, ,

Then knowing that the Perimeter is given by:

,

make the substitution:



Then solve for







For your problem: , , and provided your question actually reads (or is supposed to read): "The width is 4 inches more than the length". Do the arithmetic.

John